[Brahms-l] Shift report 20050621 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 21 2005 - 07:47:01 EDT
Shifter leader was SB
Shifters were 


<br>
<br>1. Beam from previous shift continued until 3:40 AM. C-AD has been <br>trying to get a new store but did not succed so far due to a <br>problem with RHIC CORRECTOR POWER SUPPLY 

<br>2. Runs 15703 - 15707
<br>Trig2(445k),Trig3(160k),Trig4(97k),Trig5(612k)

<br>3. No detector problems
<br>-------------------------------------------------
<br>Time             Blue Pol.(%)     Yellow Pol.(%)
<br>-------------------------------------------------
<br>1:30  AM       -46.29 +/- 2.34    -47.51 +/- 2.33
<br>At beam dump   -47.00 +/- 2.34    -42.35 +- 2.39


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 21 07:47:12 2005

This archive was generated by hypermail 2.1.8 : Tue Jun 21 2005 - 07:47:23 EDT